# Get the Agents Drunk: Memory Perturbations in Autonomous Agent-based Recommender Systems

### v1

@@
Abstract.
Large language model-based agents are increasingly used in recommender systems (Agent4RSs) to achieve personalized behavior modeling. Specifically, Agent4RSs introduces memory mechanisms that enable the agents to autonomously learn and self-evolve from real-world interactions. However, to the best of our knowledge, how robust Agent4RSs are remains unexplored. As such, in this paper, we propose the first work to attack Agent4RSs by perturbing agents’ memories, not only to uncover their limitations but also to enhance their security and robustness, ensuring the development of safer and more reliable AI agents.
Given the security and privacy concerns, it is more practical to launch attacks under a black-box setting, where the accurate knowledge of the victim models cannot be easily obtained. Moreover, the practical attacks are often stealthy to maximize the impact. To this end, we propose a novel practical attack framework named DrunkAgent. DrunkAgent consists of a generation module, a strategy module, and a surrogate module. The generation module aims to produce effective and coherent adversarial textual triggers, which can be used to achieve attack objectives such as promoting the target items. The strategy module is designed to ‘get the target agents drunk’ so that their memories cannot be effectively updated during the interaction process. As such, the triggers can play the best role. Both of the modules are optimized on the surrogate module to improve the transferability and imperceptibility of the attacks. By identifying and analyzing the vulnerabilities, our work provides critical insights that pave the way for building safer and more resilient Agent4RSs. Extensive experiments across various real-world datasets demonstrate the effectiveness of DrunkAgent.
Adversarial Attacks, Recommender Systems, Large Language Model, Generative Agents, Collaborative Filtering
copyright: noneccs: Security and privacy Web application security
1. Introduction
Recommender systems (RSs) have evolved into a vital component of the modern digital ecosystem. From e-commerce sites (e.g., Amazon, Taobao) and streaming platforms (e.g., YouTube, TikTok) to social media networks (e.g., Facebook, Instagram), RSs power personalized experience, drive user engagement of these online platforms, and further influence market dynamics through item promotions (Yang et al., 2024, 2023). The foundation of the RSs lies in modeling users’ preferences based on their historical interactions with items, such as purchases and ratings, known as collaborative filtering (CF). Despite the success of CF techniques (Koren et al., 2009; He et al., 2017, 2020), they often struggle with data sparsity and cold-start problems (Sachdeva and McAuley, 2020). To alleviate the problems, the auxiliary information with rich semantics (e.g., item descriptions and user reviews) has been integrated into RSs (Zhang et al., 2019). Given the increasing volume and complexity of recommendation data, the RSs encounter difficulties in scaling effectively and capturing textual knowledge about users and items sufficiently, due to they generally rely on traditional deep learning (DL) models like CNNs and RNNs (Zhao et al., 2023).
Figure 1. An exemplar of attacks on agent-based recommender systems, where the target item’s descriptions have been manipulated to perturb the memory of the target agent for promoting the target item to more normal users. By investigating this, our work reveals emerging vulnerabilities in memory-augmented agents and provides insights for building more secure recommender systems.
Recently, large language models (LLMs) have shown potential to mitigate the above problems and revolutionize RSs (Zhao et al., 2023), as they are primarily built upon advanced transformer-based architectures and pretrained on diverse and extensive datasets. Most of works carefully design textual prompts for LLMs based on user-item interactions (Geng et al., 2022; Bao et al., 2023; Zhang et al., 2023; Wu et al., 2024). As such, the abundant knowledge inside the LLM’s training corpora can be fully exploited. However, the LLM-based RSs (LLM4RSs) struggle to capture underlying behavioral patterns, due to the gap between universal language modeling and personalized behavior modeling (Zhang et al., 2024b). Integrating agents built upon LLMs equipped with additional modules offers a promising solution (Durante et al., 2024; Huang et al., 2024), allowing to autonomously explore and learn from external sources. Among all the added modules, memory is a key component to support agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task. More broadly, environment can be any contextual factors (e.g., item titles, user-item interactive behaviors) that influence the agent’s decisions (Zhang et al., 2024a). interactions, which plays a crucial role in helping agents accumulate historical experience, retrieve informative knowledge and self-evolve in real-world environments (Zhang et al., 2024a; Wang et al., 2024b). Consequently, the agent-based RSs with powerful LLMs (Huang et al., 2024; Zhang et al., 2024b) can also leverage memory modules that store the information explored from the external environments to further improve the recommendation quality. Fig.  1 gives an example of an item agent interacting with its situated environment (e.g., users and other items), where the learning knowledge will be placed in its memory. Despite the aforementioned success, the safety vulnerability of the generative agent-based RSs (Agent4RSs) under adversarial attacks has yet to be explored. Such attacks can gain insights into the robustness of the systems and help develop effective defensive mechanisms. As recommender agents are integrated into more complex real-world environments, understanding how these attacks might compromise their behaviors is critical for building trustworthy and robust recommender agents (Chua et al., 2024).
Due to security and privacy considerations, the access of accurate knowledge of the victim model, including architectures, gradients, parameters, etc., is generally restricted in practice. As such, practical attacks are mainly under a black-box setting (Ning et al., 2024; Yang et al., 2024; Lin et al., 2022; Zhang et al., 2024d; Zeng et al., 2023). Practical attacks are often required to have transferability (i.e., effective against different black-box RSs). Moreover, to maximize the impact of the attack and the number of affected users, practical attacks are often rewarded by imperceptibility (i.e., difficult to be discovered) (Liu and Larson, 2021).
Conventional black-box shilling attacks (Wang et al., 2024a), which inject well-crafted fake user profiles (e.g., a set of users’ numerical ratings on items) into the victim model for certain attack objectives (e.g., promote their target items for profits), have been shown to effectively poison DL-based RSs. However, most attacks have difficulty in processing textual inputs (e.g., item titles and descriptions) and lack context awareness, making it challenging to effectively target LLM-empowered RSs that rely primarily on text inputs to generate natural language responses (Ning et al., 2024). Item providers are generally responsible for creating and maintaining their own product representations on the platforms with recommendation services in practice (Yang et al., 2024; Cohen et al., 2021; Liu and Larson, 2021) (e.g., YouTube allows creators to upload item titles and Amazon enables sellers to upload item descriptions), and current RSs depend heavily on these representations (Zhang et al., 2024d; Zhao et al., 2023). As such, security vulnerabilities in these systems naturally emerge. The vulnerabilities can easily be exploited by profit-driven merchants manipulating product representations, as shown in Fig.  1. Compared to the shilling attacks that involve hiring online writers to introduce fake reviews to indirectly promote the items, this type of attack is more direct, efficient, and lower in economic cost (Zhang et al., 2024d).
Recent black-box textual attacks on LLM4RSs are far from practical. Classical text attacks that are successful in text mis-classifications, such as TextFooler (Jin et al., 2020) and PuncAttack (Formento et al., 2023), have been directly applied for LLM4RSs (Zhang et al., 2024d). However, the effectiveness of the attacks is limited, due to there is a gap between the attack goals of the text attacks and RS attacks. Trivial and ChatGPT attacks (Zhang et al., 2024d), on the other hand, are proposed to be tailored for LLM4RSs, which increases the items’ attractiveness by appending positive words to the titles. Nevertheless, such attacks are not optimized on any black-box RSs, making them poorly transferable to the Agent4RSs. Moreover, since the attacks mainly target modifications to item titles, where titles’ brevity and frequent exposure to users tend to make even small perturbations noticeable.
To tackle the above problems, in this paper, we propose a new practical attack framework named DrunkAgent. Instead of manipulating the item titles, to enhance the imperceptibility of attacks, we propose to perturb the description of the target item, where length and context can easily hide changes. DrunkAgent is composed of three modules: generation module, strategy module and surrogate module. The generation module is designed to produce effective adversarial textual description triggers, which can be used to promote the target items to as many normal users as possible on the agent-powered RSs. The triggers are coherent, semantically meaningful and human-readable text, making the attack stealthy. The strategy module aims to ‘get the target agents drunk’, that is, to confuse the agents’ memory mechanisms so that they cannot be effectively updated during agent-environment interactions. In this way, the triggers can play the best role. To avoid frequently querying the victim models to induce suspicion and mimic the black-box settings in the real world, the generation module and the strategy module are optimized on the surrogate module to improve the transferability and imperceptibility of the practical attacks.
Beyond demonstrating the attack capabilities of DrunkAgent, our broader goal is to reveal and understand the overlooked vulnerabilities in Agent4RSs. Specifically, by analyzing how agents’ memory mechanisms can be influenced through carefully crafted and stealthy textual triggers, our work provides a new lens into the emerging attack surfaces introduced by memory-augmented, LLM-based agents. We believe that this exploration can offer valuable insights to the community for designing more resilient agent architectures, guiding future efforts on defense mechanisms such as memory validation or input sanitization.
@@ 1. Introduction
  • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs. To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agent-powered RSs, aiming to inform future robustness research.
  • We propose to perturb the description of the target item for effective memory perturbations, which is an unnoticeable and crucial property of current agentic recommendations.
  • We propose a novel practical attack framework on the agent-powered RSs named DrunkAgent, which can create adversarial descriptions of target items with transferability and imperceptibility.
  • Our DrunkAgent significantly outperforms state-of-the-art black-box attacks against various LLM-based agentic RSs on real-world datasets in terms of transferability and imperceptibility. Moreover, state-of-the-art defense strategies are ineffective to DrunkAgent, demonstrating the need for improved defense mechanisms to guide more secure Agent4RSs.
2. Problem Formulation
2.1. Victim LLM-powered Agentic RSs
@@
We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively. $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile). LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (Geng et al., 2022; Zhao et al., 2023; Bao et al., 2023; Zhang et al., 2023). Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
(1) $\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$

### v2

# DrunkAgent: Stealthy Memory Corruption in LLM-Powered Recommender Agents

@@
Abstract
Large language model (LLM)-powered agents are increasingly used in recommender systems (RSs) to achieve personalized behavior modeling, where the memory mechanism plays a pivotal role in enabling the agents to autonomously explore, learn and self-evolve from real-world interactions. However, this very mechanism, serving as a contextual repository, inherently exposes an attack surface for potential adversarial manipulations. Despite its central role, the robustness of agentic RSs in the face of such threats remains largely underexplored. Previous works suffer from semantic mismatches or rely on static embeddings or pre-defined prompts, all of which hinder their applicability to systems with dynamic memory states. This challenge is exacerbated by the black-box nature of commercial RSs.
To tackle the above problems, in this paper, we present the first systematic investigation of memory-based vulnerabilities in LLM-powered recommender agents, revealing their security limitations and guiding efforts to strengthen system resilience and trustworthiness. Specifically, we propose a novel black-box attack framework named DrunkAgent. DrunkAgent crafts semantically meaningful adversarial textual triggers for target item promotions and introduces a series of strategies to maximize the trigger effect by corrupting the memory updates during the interactions. The triggers and strategies are optimized on a surrogate model, enabling DrunkAgent transferable and stealthy. Extensive experiments on real-world datasets across diverse agentic RSs, including collaborative filtering, retrieval augmentation and sequential recommendations, demonstrate the generalizability, transferability and stealthiness of DrunkAgent.
1 Introduction
The advent of large language models (LLMs) has ushered a transformative paradigm in recommender systems (RSs). Early LLM-based RSs primarily relied on verbalizing user-item interactions into prompts to guide recommendations (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; wu2024coral, ). However, such approaches often struggle to capture personalized behavioral patterns, due to a fundamental gap between interaction dynamics and generic language modeling (zhang2024agentcf, ). To bridge this gap, recent works have introduced autonomous agents built upon LLMs and augmented with modules, such as memory, planning, and tool-use, laying the foundation for agentic RSs (durante2024agent, ; huang2024foundation, ). As personalized user experiences become central to digital platforms, LLM-powered agentic RSs are rapidly emerging as the next evolutionary step (huang2025towards, ), yet their robustness, particularly against adversarial threats, remains largely underexplored.
At the core of agentic systems lies the memory module, which plays a pivotal role in supporting agent-environment [^1] In a narrow sense, environment is the object that the agent needs to interact with to accomplish the task. More broadly, environment can be any contextual factors that influence the agent’s decisions (zhang2024survey, ). interactions (zhang2024survey, ; wang2024survey, ). Specifically, the module retains historical interactions, allowing agents to accumulate contextual knowledge, adapt to evolving user preferences, and refine their behavior over time through interaction-driven updates (huang2025survey, ), as shown in Fig. 1. However, this very mechanism that empowers agent autonomy also introduces a new and overlooked attack surface.
Adversarial attacks targeting memory often leave a lasting impact. Given that attacks are generally persistent rather than one-off events, they can be readily exploited in non-stationary environments to take advantage of the agent’s continual learning process (huang2025towards, ; zhang2024agentcf, ), leading to repeated integration of adversarial signals into its evolving internal state. This results in long-term memory contamination and hence strengthens a drift in user preferences. Under sequential and retrieval modeling (huang2024foundation, ; huang2025survey, ), these perturbations persist longer by influencing the agent’s ongoing behavior trajectory. Perturbed memory traces can gradually bias the agent’s belief state through temporal dependencies, while also being frequently retrieved in future recommendations. This dual exposure amplifies the longevity and impact of even small-scale attacks. In this work, we take the first step toward enhancing the robustness and security of the LLM-powered agentic RSs by systematically investigating their adversarial vulnerabilities, with a particular focus on stealthy memory corruption attacks.
Figure 1: In LLM-powered agentic RSs, agents autonomously interact with their situated environments including other agents to collaboratively optimize their memories over time (Left), with recommendations being generated based on these memories (Right). DrunkAgent perturbs the memory of the target item agent, where the attack strategy aims to ‘get the target agent drunk’ so that the adversarial triggers can be injected into the target memory (The Temporal Snapshot of the Left). As a result, the target item can be promoted to more recommendation lists of normal users (Right).
Existing works on attacks against RSs falls into three largely independent threads: 1) poisoning attacks on traditional collaborative filtering (CF) models (wang2024poisoning, ), which inject fake user profiles to corrupt embeddings; 2) adversarial text attacks on LLM-based recommenders (zhang2024stealthy, ; ning2024cheatagent, ), perturbing input prompts to misguide outputs; and 3) agentic RS architectures that enhance personalization through memory-augmented LLMs (huang2025survey, ). However, these threads remain disconnected. Poisoning methods often assume static embeddings and focus on numerical perturbations, creating a semantic gap with text-driven systems. Textual attacks tend to focus narrowly on pre-defined prompts, limiting their applicability in agentic RSs with dynamic memory updates. Meanwhile, the agentic studies neglect security analyses, leaving critical vulnerabilities unexplored. This challenge is exacerbated by the black-box nature of commercial RSs, where attackers lack access to model parameters and training pipelines. Moreover, a critical threat vector aligned with real-world platform abuse incentives lies in item promotion attacks. Poisoning attacks rely on fake user injections, making promoting items costly and indirect, while textual attacks typically focus on degrading overall performance or causing mis-classification, overlooking practical promotion-driven threats (wang2024poisoning, ; ning2024cheatagent, ; zhang2024stealthy, ).
To tackle the above problems, in this paper, we propose DrunkAgent, the first black-box framework to exploit memory-based vulnerabilities in LLM-powered recommender agents. DrunkAgent operates under two principles: 1) memory confusion, where adversarial textual inputs to disrupt agents’ ability to retain and update interaction histories, leading to persistent distortions in their memories and hence shifts in user preferences, and 2) semantic stealth, where low-perplexity adversarial text are crafted to appear linguistically natural and coherent, enabling undetected memory corruption. Specifically, DrunkAgent crafts semantically meaningful adversarial textual triggers for target item promotions and adopts a series of strategies to corrupt the memory updates of the target item agents during the environmental interaction process, allowing the triggers to achieve maximum impact, as shown in Fig. 1. To mitigate the risk of raising suspicion from frequent queries on victim models in prior methods (zhang2024stealthy, ; ning2024cheatagent, ; wang2024poisoning, ) and to adhere to realistic black-box constraints, DrunkAgent introduces a surrogate model to optimize the triggers and strategies to further improve the attack transferability and stealthiness. To comprehensively evaluate DrunkAgent, we conduct extensive experiments across real-world datasets from different domains, a variety of attack baselines with varying perturbation strengths, representative black-box agentic RSs featuring diverse designs and recommendation tasks, including collaborative filtering, retrieval augmentation, and sequential recommendations, widely adopted stealthiness evaluation methods, and cutting-edge defense mechanisms.
@@ 1 Introduction
  • We identify a safety vulnerability in autonomous agent-based RSs with powerful LLMs. To the best of our knowledge, we present the first formal and systematic study of adversarial textual attacks on the agentic RSs, aiming to inform future robustness research.
  • We propose DrunkAgent, a novel black-box attack framework specifically designed for the agent-powered RSs, which crafts adversarial triggers and tailored strategies for effective memory perturbations of target item agents.
  • DrunkAgent consistently outperforms state-of-the-art black-box attacks against a wide range of agentic RSs and under multiple stealthiness evaluation methods on real-world datasets, in terms of both transferability and stealthiness. Moreover, it remains resistant to existing advanced defense mechanisms, exposing critical blind spots in current countermeasures and underscoring the urgent need for adaptive defenses.
2 Problem Formulation
2.1 Victim LLM-powered Agentic RSs
@@
We use $\mathcal{Y}=\{y_{u,v}:u\in\mathcal{U},v\in\mathcal{V}\}$ to denote the records of the user-item interaction matrix in the recommendation space, where $\mathcal{U}$ and $\mathcal{V}$ represent the set of real users and the item universe, respectively. $\mathcal{V}_{u}=\{v\in\mathcal{V}:y_{u,v}\neq 0\}$ indicates the set of items that have been interacted by $u$ (i.e., the user’s profile). LLM-powered RSs generally conduct recommendations by integrating user historical behavior sequences and/or item and/or user features with personalized prompt templates (geng2022recommendation, ; bao2023tallrec, ; zhang2023collm, ; zhao2023recommender, ). Let $\mathcal{X}_{u,v}$ represent the prompts of LLM-powered RSs (denoted as $\text{LLM}_{\Theta}$ with parameters $\Theta$), the preference function can be formulated as below.
$\mathcal{R}_{u}=f_{\text{LLM}_{\Theta}}(\mathcal{X}_{u,v}),$ (1)